find an equation of the line having the given slope and containing the given point. m=3,(4,1).
The equation of the line in slope-intercept form is y=??

Answer by ewatrrr(24785)


Hi
line having the given slope and containing the given point. m = 3, (4,1).


y = 3x + b |using ordered pair Pt(4,1) to solve for b
1 = 12 + b
-11 = b
y = 3x -11